### 13 Mar 2024 v2

# Mipha: A Comprehensive Overhaul of Multimodal Assistant with Small Language Models

Abstract
Multimodal Large Language Models (MLLMs) have showcased impressive skills in tasks related to visual understanding and reasoning. Yet, their widespread application faces obstacles due to the high computational demands during both the training and inference phases, restricting their use to a limited audience within the research and user communities. In this paper, we investigate the design aspects of Multimodal Small Language Models (MSLMs) and propose an efficient multimodal assistant named Mipha, which is designed to create synergy among various aspects: visual representation, language models, and optimization strategies. We show that without increasing the volume of training data, our Mipha-3B outperforms the state-of-the-art large MLLMs, especially LLaVA-1.5-13B, on multiple benchmarks. Through detailed discussion, we provide insights and guidelines for developing strong MSLMs that rival the capabilities of MLLMs. Our code is available at https://github.com/zhuyiche/Mipha.
1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) have demonstrated exceptional visual understanding and reasoning performances across a range of tasks such as visual question answering Team et al. (2023); Alayrac et al. (2022); Wang et al. (2023) and referring comprehension Achiam et al. (2023). Benefiting from the scaling law of Transformer architecture Vaswani et al. (2017) and web-scale training data sources, these models have become foundational in the field of artificial intelligence, with their parameters increasing from billions to trillions Achiam et al. (2023); Touvron et al. (2023b); Chowdhery et al. (2023). However, the deployment of these models is often hindered by their substantial computational costs and memory consumption in both the training and inference phases, which limits their popularization across the broader research and user communities.
How can we boost the inference speed of MLLMs? Delving into the computational costs of MLLMs, it becomes evident that the Large Language Models (LLMs), primarily tasked with transforming visual-text pairs into textual outputs, account for a substantial computational load. In MLLMs, the visual encoder, such as CLIP-ViT-L, typically only comprises around 0.4 billion parameters, while the language model within MLLMs can escalate to between 7 and 65 billion parameters.
Intuitively, reducing the computational demands of the language model could lead to a significant decrease in overall inference costs. The rapid advancements by the open-source community in developing Small Language Models (SLMs), with parameter sizes ranging from 1 billion to 3 billion (i.e., Phi Gunasekar et al. (2023); Li et al. (2023e); Javaheripi et al. (2023)), enable us to construct vision-language models more compact than traditional MLLMs. However, as the parameter count of these smaller language models is reduced, their capabilities often do not measure up to those of their larger counterparts, potentially diminishing the effectiveness of small multimodal models. This motivation drives us to explore the development of multimodal assistants using small language models that can deliver competitive performance comparable to their counterparts Dai et al. (2023b); Bai et al. (2023); Ye et al. (2023a, b); Liu et al. (2023a), which rely on large language models.
Figure 1: Evaluation Benchmarks Overview. We evaluate our model variants on academic-task-oriented benchmarks (left) as well as instruction-following benchmarks (middle). The answer is predicted by our proposed efficient MSLM: Mipha-3B. Additionally, we explore three key design spaces of MSLMs: 1) visual representation, 2) language model, and 3) optimization strategy (right).
In this paper, we present an in-depth analysis of the impact of various design elements in MSLMs on their downstream performance. A brief summary of the design space that we explored is illustrated in Fig 1 (right). Our investigation dissects the existing MSLMs into three key components: the language model, the visual representation, and the optimization strategy. Our thorough empirical research leads us to several findings that diverge from the conventional wisdom established by prior studies on Multimodal Large Language Models. For example:
  • Increasing image resolution is not a silver bullet. In some benchmarks, images with a resolution of 224 pixels outperform those with 448 pixels.
  • Contrary to earlier findings that suggest finetuning the visual backbone negatively impacts MLLMs, we underscore the importance of simultaneously finetuning both the visual backbone and the language model for MSLMs.
  • While the prevailing trend among MLLMs is to employ instruct-tuned language models, such as Vicuna Chiang et al. (2023), our analysis reveals that instruction tuning—be it through supervised fine-tuning or reinforcement learning from human feedback (RLHF)—is not essential.
Building on these findings, we present a new family of MSLMs named Mipha, scaled from 1.7B to 3B, which surpasses the performance of leading open-source MLLMs including LLaVA-1.5 and Qwen-VL, without the need for additional training data. Remarkably, our Mipha-3B models demonstrate superior performance across a majority of benchmarks when compared to the 7B MLLM variants, such as LLaVA-1.5, Qwen-VL, and InstructBLIP, and in some cases, even outperform 13B MLLM counterparts. Overall, we believe that our analysis provides a practical new perspective on optimizing effective MSLMs. Overall, our contributions are the follows:
  1. 1.
    We offer a detailed examination of the design landscape for Multimodal Small Language Models, examining visual components, language models, and optimization strategies from a comprehensive perspective. Our empirical research provides unique findings, presenting multiple conclusions that contrast with prior studies on Multimodal Large Language Models.
  2. 2.
    By integrating our empirical discoveries, we unveil a series of Multimodal Small Language Models named Mipha. This advancement underscores the practical impact of our findings.
  3. 3.
    Demonstrating the effectiveness of our approach, we show that our flagship model, Mipha-3B, outperforms existing 7B and 13B models on various benchmarks without requiring additional data for training. This marks a significant advancement for multimodal models with fewer than 3 billion parameters, establishing a new benchmark in the field.
2 Related Work
Language Models. Arming with scaled-up data and models, Large Language Models (LLMs) have demonstrated emergent capabilities like zero-shot generalization and in-context learning ability Chowdhery et al. (2023); Chung et al. (2022); Achiam et al. (2023); Touvron et al. (2023a, b). Despite these advances, LLMs are computationally intensive, demanding significant computational resources for training and inference. This backdrop has inspired the development of smaller language models Gunasekar et al. (2023); Li et al. (2023e); Javaheripi et al. (2023); Biderman et al. (2023); Liu et al. (2024); Zhang et al. (2024); Stability (2024); Google (2024), which, despite having fewer than 3 billion parameters, exhibit strong generalization capabilities in language understanding and question-answering tasks, rivaling their larger counterparts such as 7 billion parameter models. In our research, we explore the integration of these compact language models with visually-conditioned models, aiming to empower neural networks with the dual ability to comprehend and interpret both textual and visual information.
Multimodal Large Language Models. Multimodal Large Language Models (MLLMs) connect vision and language and extend the reasoning ability of LLMs to process with multimodal input. Numerous works have been proposed in this direction Li et al. (2023c); Dai et al. (2023b); Zhu et al. (2024a); Liu et al. (2023b); Team et al. (2023); Chen et al. (2023a); Li et al. (2023a); Lin et al. (2023); Bavishi et al. (2023), which most works differ based on their adapter design, training strategy, instruction-tuning/pretraining datasets. Leveraging their complex reasoning and advanced understanding abilities, LLMs serve as the "brain" of these systems, interpreting visual elements based on text queries and generating coherent, fluent textual responses in dialogues. Currently, the bulk of the computational demand stems from this "brain", with parameter counts ranging from 7 billion to 65 billion for publicly available models. This substantial requirement for computational resources presents a significant challenge for deployment on mobile or edge devices, making the inference cost-prohibitive and limiting the accessibility of MLLMs to a wider audience. This work aims to explore MSLMs that can achieve performance competing with MLLMs over 7B parameters.
Multimodal Small Language Models. Recently, a select number of studies Shao et al. (2024); He et al. (2024); Chu et al. (2024); Zhu et al. (2024b); Lin et al. (2024); Wei et al. (2024); Zhang et al. (2022) have delved into the exploration of Multimodal Small Language Models (MSLMs) from diverse angles. Several works, i.e, LLaVA-Phi Zhu et al. (2024b), utilize pretrained small language models as the core of multimodal models to reduce computational demands. The MobileVLM Chu et al. (2023, 2024) series focuses on the innovative design of projectors to enhance the performance of MSLMs. MiniCPM-V min (2024) seeks to refine the training approach, and Bunny He et al. (2024) examines the effects of augmenting the volume of training data. Our paper put a thorough investigation into the design space for training an effective MSLM.
3 Preliminaries
In this section, we provide an overview of model architecture and evaluation benchmarks.
3.1 Overview of Model Architecture
The architecture of mainstream MLLMs Liu et al. (2023a); Zhu et al. (2024a); Ye et al. (2023b) typically comprises three primary components: 1) a visual representation backbone $V_{\psi}$, 2) a vision-language projector $F_{\omega}$ aimed at aligning visual and textual domains, and 3) a language model $LM_{\theta}$.
Visual Representation Backbone. Given an input image $X_{img}$, we consider a visual representation backbone to output the visual features $V_{img}\in\mathbb{R}^{L\times h_{v}}$, where $V_{img}=V_{\psi}(X_{img})$. The $L$ denotes the number of patches. All of our experiments consider only the grid features before and after the last transformer layer.
Vision-Language Projector. Then, a vision-language projector is employed to convert the visual features $V_{img}$ to the text embedding space $T_{img}\in\mathbb{R}^{L\times h_{t}}$, where $T_{img}=F_{\omega}(V_{img})$. Following LLaVA-1.5 Liu et al. (2023a), we employ a two-layer MLP as our projector across all of our model variants, which can effectively align visual and textual features.
Language Model (LM). Finally, based on $T_{img}$ and the question embedding $E_{t}$, a language model is used to give the answer $A$, where $A=LM_{\theta}(text{cat}(T_{img},E_{t}))$. While previous studies mainly focus on language models with more than 7 billion parameters, such as Vicuna or LLaMA, our work shifted attention to multimodal models that are equipped with SLMs.
Figure 2: Selection of Small Language Models. We evaluate four open-sourced SLMs on 8 benchmarks, and Phi-family showcases the best performance (left &\&& middle). We discovered that MSLMs equipped with Phi-2-2.7B is able to identify that the monkey in the image is performing, a subtlety that other models fail to recognize (right).
Table 1: Comparison of small language models on Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2021), BIG-Bench Hard (BBH) Srivastava et al. (2022), GSM8K Cobbe et al. (2021), and HumanEval Chen et al. (2021).
Method MMLU BBH GSM8K HumanEval
Language Understanding Instruction Following Math Coding
Pythia-1.4B 25.4 - - 4.3
StableLM-1.6B 45.9 37.7 52.5 35.4
Qwen1.5-1.8B 46.8 24.2 38.4 20.1
Phi-1.5-1.3B 43.9 - 44.6 41.4
Pythia-2.8B 28.3 - - 5.1
Phi-2-2.7B 56.7 43.4 57.2 47.6
3.2 Evaluation Benchmarks.
We conduct empirical studies on a collection of both academic-task-oriented benchmarks and recent benchmarks specifically proposed for instruction-following MLLMs, totaling 8 benchmarks. For academic-task-oriented benchmarks, VQA-v2 Goyal et al. (2017) and GQA Hudson and Manning (2019) evaluate the model’s visual perception capabilities on open-ended short answers. ScienceQA Lu et al. (2022) is used to evaluate the zero-shot generalization on scientific question answering. TextVQA Singh et al. (2019) contains text-rich visual question answering. We also employ recent benchmarks proposed for instruction-following MLLMs. POPE Li et al. (2023d) evaluates the model’s degree of hallucination on three sampled subsets of COCO Lin et al. (2014). MME Fu et al. (2023) Benchmark evaluates MLLMs’ perception and cognition capabilities, and MMBench Liu et al. (2023d) evaluates the model’s answer robustness with all-around shuffling on multiple-choice answers. MM-Vet Yu et al. (2023) evaluates the model’s capabilities in engaging in visual conversations on a diverse range of tasks. Some examples are illustrated in Fig 1.
4 Exploring Crucial Aspects on Efficient MSLMs
There are a few prior studies that have explored different aspects of enhancing MLLMs Karamcheti et al. (2024); Lu et al. (2023). However, these works primarily focus on multimodal assistants powered by large language models. Intuitively, the model insights obtained from multimodal large language model should also apply to multimodal smaller language models. Nevertheless, in the following section, the contrary: techniques effective for large-scale models do not always seamlessly transfer to small-scale models. We also demonstrate several new observations that can improve the performance of MSLMs. We highlight that all our experiments in this section utilize the same amount of training data, and unless otherwise stated, all training components are maintained consistently for a fair comparison.
4.1 What makes Small Language Model good for Multimodal Assistant?
In this section, we delve into the methodology for selecting and training an effective Small Language Model (SLM) that enhances the performance of Multimodal Small Language Models (MSLMs). Our discussion is twofold. Initially, we evaluate various SLMs to determine which model exhibits optimal performance upon integration with a visual backbone. Subsequently, we investigate the MSLMs’ performance with base or instruct-tuned language models.
4.1.1 Which Small Language Model should we adopt?
As the "brain" of multimodal models, the selection of pretrained small language models is crucial for crafting high-performing Multimodal Small Language Models (MSLMs). We evaluate a spectrum of recently introduced Small Language Models (SLMs) with sizes ranging from 1.3B to 2.8B parameters. This includes models such as Phi-2 (2.7B) Javaheripi et al. (2023), Phi-1.5 (1.3B) Li et al. (2023e), Qwen-1.5 (1.8B) Bai et al. (2023), Pythia (available in both 1.4B and 2.8B sizes) Biderman et al. (2023), and StableLM (1.6B) Stability (2024). For analytical clarity, we classify these SLMs into two categories: those models possess over 2.7B parameters, and the remainder, feature less than 2B parameters.
Figure 3: Base vs. Instruct-tuned LMs. For MLLMs, we explore the differences using base LM and instruct-tuned LM. While the quantitative performance metrics appear similar, the qualitative results reveal differences (left). For example, when comparing the responses generated by MLLMs equipped with either a Base or SFT LM to those from models finetuned with RLHF or Safe-RLHF, the latter is observed to be more verbalized (right).
As illustrated in Figure 2 (left), we assess the performance of MSLMs that incorporate language models with fewer than 2 billion parameters. Phi-1.5-1.3B, despite having fewer parameters compared to its counterparts, still manages to display competitive or even superior performance. To delve into the reasons behind Phi-1.5’s outperformance, we benchmark these SLMs against four tests traditionally used for LLMs, as depicted in Table 1. These benchmarks include the MMLU for language understanding, BBH for instruction following, GSM8K for mathematics, and HumanEval for coding. Our observations reveal that Phi-1.5-1.3B remains competitive or outperforms models with a larger parameter count, such as Pythia-1.4B, StableLM-1.6B, and Qwen-1.8B. This suggests that Phi-1.5-1.3B makes efficient use of its parameters, particularly in the areas of reasoning and language understanding that these benchmarks measure, enhancing its transferability to multimodal contexts. Our findings indicate that a strong small language model can lead to better multimodal performance.
Moreover, when comparing Phi-2-2.7B with Pythia-2.8B, two relatively larger SLMs, the superiority of Phi-2 is evident. Aside from POPE, where models with Pythia show slightly fewer object hallucinations, Phi-2 models achieve markedly better outcomes on other metrics. This aligns with expectations, as Phi-2 was released after Pythia and was pretrained on a considerably larger dataset—1.4 trillion tokens versus 300 billion tokens.
Additionally, the qualitative results corroborate our findings. MSLMs based on the Phi models provide detailed descriptions derived from the images. Notably, the MSLM equipped with Phi-2-2.7B goes a step further by recognizing that the monkey is performing, which showcases a robust capacity for scene reasoning. Additional examples and detailed analyses are provided in the appendix.
Base vs. Instruct-tuned LMs. The most state-of-the-art MLLMs Liu et al. (2023a); Dai et al. (2023b) leverage instruct-tuned LM, such as Vicuna, as the default backbone. However, employing instruction tuning on LMs means extra training time, and it also brings bias and regressions in performance Ouyang et al. (2022). As such, we investigate whether instruction tuning is necessary for MSLMs. We consider the following four settings:
  • Base. The base language model (LM) inherits the weights from pretrained models as is. Typically, this base LM exhibits limited proficiency in following instructions, due to its training on a general corpus without specific emphasis on instruction adherence.
  • Supervised Fine-tuning (SFT). Supervised Fine-Tuning (SFT) refines a pretrained model’s parameters using a targeted, often smaller, dataset with labeled examples to enhance its performance for a specific task. This technique builds upon the broad capabilities acquired during the model’s initial training phase, sharpening them to boost accuracy, relevance, or suitability for a designated domain or application. Within the realm of Multimodal Large Language Models (MLLMs), employing a language model enhanced through SFT, such as Vicuna, serves as a prevalent strategy.
  • Reinforcement Learning from Human Feedback (RLHF). RLHF Ouyang et al. (2022) finetunes LMs by integrating preferences or feedback directly from humans. This method refines the capabilities of base LMs to function effectively as dialogue agents. This approach not only enhances the responsiveness and relevance of the LMs to user queries but also aligns their outputs more closely with human values and expectations, thereby broadening their utility as interactive tools.
  • Safe Reinforcement Learning from Human Feedback (Safe-RLHF). Safety alignment is a critical step to mitigate the risk of LMs generating undesirable outputs in response to certain prompts Liu et al. (2023c). In our approach, we employ Safe-RLHF Dai et al. (2023a), a variant of RLHF, which incorporates safety preferences throughout the RLHF process. This method is designed to ensure that the model’s responses adhere to safety guidelines, effectively preventing inappropriate outputs.
The experimental findings are presented in Fig 3 (left), revealing that there is no universal solution to enhancing model generalization. The comparative analysis across all 8 benchmarks shows minimal variance among the four configurations tested. Notably, the MSLM leveraging the base Phi-2 model excels in the SQA-I and MM-Vet benchmarks. This superiority can likely be attributed to Phi-2’s extensive pretraining on large code datasets, which enhances its ability to generalize in science question-answering contexts and generate open-ended responses. While the performance differences among instruct-tuned models are modest, RLHF emerges as the most versatile performer across the eight benchmarks. Despite the close numerical results, qualitative analysis reveals a distinct divergence between instruct-tuned and base models. As depicted in Fig 3 (right), when tasked with describing an artwork image, the responses from models with base LM and SFT LM adopt a formal tone, employing terms like "religious elements" and "three-dimensional sculptures", reminiscent of textbook descriptions. Conversely, models finetuned with RLHF and Safe-RLHF LMs offer descriptions in a more human-like, conversational manner, akin to a museum commentary.
Takeaway. The benefit of employing an instruct-tuned LM for MSLMs over a base LM is not necessarily clear-cut, irrespective of the instruction tuning approach used, whether it be supervised fine-tuning, reinforcement learning from human feedback (RLHF), or safety-aware RLHF. This finding contrasts with prior research, which indicated that using an instruct-tune language model like Vicuna offers advantages over a base language model like LLaMA. However, qualitative differences become apparent: MSLMs developed using RLHF-based instruction tuning methods tend to be more verbose and produce responses that are more akin to human-like answers.
Figure 4: Choosing a Pretrained Vision Representation & Scaling Image Resolution. We evaluate various visual backbones, such as CLIP, SigLIP, DINOv2, and ViT-IN21K (left). We analyze model performance in relation to increasing image resolution (middle). We provide qualitative examples from multimodal models employing visual backbones at different image resolutions (right).
4.2 Visual Representations in MSLM
Pretrained Visual Representation. In multimodal models, a common approach is to employ a pretrained visual backbone that converts image features into latent embeddings. The quality of these pretrained embeddings is pivotal to the success of multimodal learning tasks. We evaluated the efficacy of four different pretrained visual backbones: 1) CLIP Radford et al. (2021), trained on a large proprietary dataset by OpenAI, 2) SigLIP Zhai et al. (2023), pretrained on the WebLI dataset, 3) ViT with IN21K, trained on the ImageNet21K Ridnik et al. (2021) dataset using vision transformer Dosovitskiy et al. (2021), and 4) DinoV2 Oquab et al. (2023), pretrained on the LVD-142M dataset. For an equitable comparison, we standardized the use of ViT-Large with a 14-patch size and an input resolution of 336 pixels. As illustrated in Figure 4 (left), it becomes evident that training on extensive visual corpora—akin to CLIP’s methodology—confers an advantage in terms of enhanced visual understanding capabilities.
Takeaway. The overarching conclusion from our experiments in this section resonates with insights gained from the training of large vision-language models. In particular, we find that the vision backbone, trained using vision-language contrastive loss (i.e., CLIP and SigLIP), outperforms other models. Interestingly, our experiments further reveal that SigLIP demonstrates a surprisingly significant performance enhancement over CLIP, despite being supervised and operating on a considerably smaller training dataset. This observation suggests that refining training methodologies for CLIP-based models could be more impactful for small Vision-Language Models (VLMs) than for their larger counterparts, indicating a potential shift in optimization strategies for enhancing model effectiveness across different scales.
Scaling Image Resolution. Image resolution plays a vital role among various design choices in neural network architecture, with past research Lu et al. (2023) indicating that, for MLLMs, increasing image resolution is more effective than scaling other dimensions, such as model width and depth. This assertion is supported by findings Karamcheti et al. (2024) that enhancing image resolution uniformly improves model performance across all benchmarks. Naturally, one might assume that this principle should apply equally to MSLMs. However, is this assumption truly valid?
We present experiments using a popular visual backbone, CLIP, with varying image sizes. The experimental results are presented in Fig 4 (middle). Contrary to the common belief that increasing image resolution is always beneficial to model generalization, our empirical results indicate that higher image resolution does not guarantee improved performance for MSLMs. For instance, a CLIP model with a 224px resolution outperformed its higher-resolution counterparts in specific benchmarks, while a 336px resolution CLIP model excelled in others.
Despite these varied results, it’s acknowledged that increased resolution can benefit tasks requiring detailed visual understanding, such as OCR, where higher resolution inputs enable more accurate recognition of details like price tags in our qualitative example in Fig 4 (right).
Takeaway. Scaling up image resolution is not a silver bullet. Although higher image resolutions offer more detailed visual inputs, they do not invariably enhance the performance of MSLMs. It is essential to find the right equilibrium between resolution and computational efficiency, tailored to the model’s application context and the need for consistent understanding and output. Efforts might be better allocated to scaling other aspects of the visual backbone, such as its depth, rather than relying solely on increased image resolution.
4.3 Optimization
Figure 5: Frozen vs. Finetuning & Full-parameter tuning vs. LoRA. We study the impact of optimization strategies on the performance of small multi-modal models. Specifically, we explored the effects of activating or freezing the visual representation backbone and LM during the instruction tuning phase on model performance (left). Additionally, we confirmed that compared to full-parameter tuning, applying LoRA to MSLMs is equally effective and can significantly alleviate the training burden (right).
Optimization strategies play a pivotal role in shaping the generalizability of a model. Most MLLMs Liu et al. (2023b, a); Zhu et al. (2024a); Ye et al. (2023b) typically adopt a two-stage optimization pipeline. The first stage, or the pretraining stage, is dedicated to harmonizing vision and language features, establishing a solid foundation for multimodal understanding. The subsequent stage focuses on instructing the model to follow users’ instructions. This section delves into the training paradigms of MLLMs, exploring the necessity of finetuning the language model and visual backbones. Additionally, we evaluate the feasibility and effectiveness of parameter-efficient fine-tuning methods, such as LoRA, in this context.
Frozen vs. Finetuning. This section examines whether the visual representation backbone and the language model should be configured as trainable or frozen. We present and analyze four different configurations that were explored.
Table 2: Multi-modal evaluation on academic-task-oriented benchmarks. VQAv2 Goyal et al. (2017); GQA Hudson and Manning (2019); VizWiz Gurari et al. (2018); $\text{SQA}^{\text{I}}$: ScienceQA-IMG Lu et al. (2022); $\text{VQA}^{\text{T}}$: TextVQA Singh et al. (2019). $*{}^{*}$he training images/annotations of the datasets are observed during training. ${}^{\dagger}$ Includes in-house data that is not publicly accessible. Column Res. is the image resolution of visual backbone. Columns PT and IT are the data sizes in the pretraining stage and the visual instruction tuning stage, respectively.
[…]
Multimodal Large Language Models
BLIP-2 Li et al. (2023c) Vicuna (13B) 224 129M - 65.0 41.0 19.6 61.0 42.5
InstructBLIP Dai et al. (2023b) Vicuna (7B) 224 129M 1.2M - 49.2 34.5 60.5 50.1
InstructBLIP Dai et al. (2023b) Vicuna (13B) 224 129M 1.2M - 49.5 33.4 63.1 50.7
Shikra Chen et al. (2023b) Vicuna (13B) 224 600K 5.5M $*{}^{*}$
IDEFICS-9B Laurençon et al. (2023) LLaMA (7B) 224 353M 1M 50.9 38.4 35.5 - 25.9
IDEFICS-80B Laurençon et al. (2023) LLaMA (65B) 224 353M 1M 60.0 45.2 36.0 - 30.9
Qwen-VL Bai et al. (2023) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 35.2 67.1 $*{}^{*}$
Qwen-VL-Chat Bai et al. (2023) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 38.9 68.2 $*{}^{*}$
mPLUG-Owl2 Ye et al. (2023b) LLaMA (7B) 448 400M 1.23M 79.4 56.1 54.5 68.7 58.2
LLaVA-1.5 Liu et al. (2023a) Vicuna (7B) 336 558K 665K $*{}^{*}$ $*{}^{*}$ 50.0 66.8 58.2
LLaVA-1.5 Liu et al. (2023a) Vicuna (13B) 336 558K 665K $*{}^{*}$ $*{}^{*}$ 53.6 71.6 61.3
Multimodal Small Language Models
MobileVLM Chu et al. (2023) M-LLaMA (2.7B) 336 558K 158K - 59.0 - 61.2 47.5
MobileVLM-v2 Chu et al. (2024) M-LLaMA (2.7B) 336 1.2M 2.4M - 61.1 - 70.0 57.5
MoE-LLaVA-3.6B Lin et al. (2024) Phi-2 (2.7B) 384 558k 1.59M $*{}^{*}$ $*{}^{*}$ 43.7 70.3 57.0
Bunny-3B He et al. (2024) Phi-2 (2.7B) 384 2M 695K 79.8 62.5 - 70.9 -
Mipha-3B (Ours) Phi-2 (2.7B) 384 558K 665K $*{}^{*}$ $*{}^{*}$ 47.5 70.9 56.6
  • Frozen ViT+SLM: In this configuration, the parameters of both the ViT and SLM are held constant, allowing the training to focus exclusively on aligning these two components via the projector.
  • Frozen SLM + Finetune ViT: Here, we freeze the small language model’s parameters and solely fine-tune the Vision Transformer.
  • Finetune LM + Frozen ViT: This strategy involves freezing the ViT while training both the projector and the small language model, an approach exemplified by LLaVA-1.5 Liu et al. (2023a).
  • Finetune ViT + SLM: This comprehensive approach activates all parameters within the MSLMs, permitting the model to simultaneously refine its visual processing capabilities and improve instruction-following proficiencies.
Our exploration into these strategies involves selectively activating or freezing the visual and language model during the instruction-tuning stage. As depicted in Fig 5 (left), our findings suggest that freezing the language model can detrimentally affect the MSLMs’ performance across various benchmarks. Furthermore, we observe that fully engaging all parameters of the MSLMs can lead to improvements on certain benchmarks.
Takeaway. Contrary to previous research that suggested fine-tuning the vision backbone could lead to a significant decrease in performance across all benchmarks Karamcheti et al. (2024), our study indicates that fine-tuning the vision encoder can actually enhance the performance of MSLMs across all evaluated benchmarks. We believe the underlying reason for this discrepancy is that SLMs possess fewer active neurons than LLMs. As a result, the visual representations in MSLMs must be more expressive than those in MLLMs to enable the language model to effectively comprehend image contents. Furthermore, our observations reveal that fine-tuning the language model is even more critical for achieving a successful MSLM implementation. We attribute this necessity to the fact that SLMs have less pre-trained data and fewer parameters, making the learning process to understand images with the assistance of visual backbones a pivotal challenge.
Full-Parameter Tuning vs. LoRA. The aforementioned section shows that fine-tuning visual backbone and LMs are necessary for MSLMs. However, full-parameter fine-tuning entails a significant increase in training costs. We investigate the potential of parameter-efficient fine-tuning techniques as alternatives to conventional full-parameter tuning. In Fig 5 (middle), we juxtapose the performance of models utilizing Low-Rank Adaptation (LoRA) against those subjected to full-parameter tuning. For the LoRA setup, we configure $r$ to be 128 and $\alpha$ to be 256. The experimental outcomes suggest that LoRA attains performance on par with its full-parameter counterparts across all evaluated benchmarks, positioning it as a viable and efficient alternative to full-parameter fine-tuning.
5 Experiments
5.1 Implementation Details
Integration of Key Design Elements. Through our empirical study, we have distilled a set of insights that collectively enhance both the downstream performance and the efficiency of models:
  • Visual Representations. We have chosen the SigLIP backbone with a resolution of 384px as our standard for visual processing.
  • Language Models. We use Supervised finetuned Phi-2-2.7B as our language backbone for Mipha-3B.
  • Optimization. Recognizing the importance of fine-tuning both the visual and language components, we undertake fine-tuning of both models using their full parameter sets during the training phase
Leveraging these insights, we developed a novel suite of models, named Mipha, scaled to 3 and 1.7 billion parameters, aiming to set new benchmarks in multimodal models. More detailed training settings are presented in the appendix.
5.2 Comparison with SOTA Models
Table 3: Multi-modal evaluation on instruction following benchmarks. POPE Li et al. (2023d); MME-P: MME Perception Fu et al. (2023); MME-C: MME Cognition Fu et al. (2023); MMB: MMBench Liu et al. (2023d); SEED: SEED-Bench Li et al. (2023b); MM-Vet Yu et al. (2023). ${}^{\dagger}$ Includes in-house data that is not publicly accessible. Columns PT and IT are the data sizes in the pretraining stage and the visual instruction tuning stage, respectively.
Method PT IT POPE MME- MMB SEED MM-Vet
P C en cn img
Multimodal Large Language Models
BLIP2-14B Li et al. (2023c) 129M - 85.3 1293.8 290.0 - - 49.7 22.4
InstructBLIP-8B Dai et al. (2023b) 129M 1.2M - - - 36 23.7 58.8 26.2
InstructBLIP-14B Dai et al. (2023b) 129M 1.2M 78.9 1212.8 291.8 - - - 25.6
Shikra-13B Chen et al. (2023b) 600K 5.5M - - - 58.8 - - -
IDEFICS-80B Laurençon et al. (2023) 353M 1M - - - 54.5 38.1 53.2 -
Qwen-VL Bai et al. (2023) ${}^{\dagger}$ ${}^{\dagger}$ 38.2 7.4 62.3 -
Qwen-VL-Chat Bai et al. (2023) ${}^{\dagger}$ ${}^{\dagger}$ 1487.5 360.71 60.6 56.7 65.4 -
mPLUG-Owl2-7B Ye et al. (2023b) 400M 1.23M 85.8 1450.2 313.2 64.5 - - 36.2
LLaVA-7B Liu et al. (2023b) 595K 158K 72.9 809.6 247.9 38.7 36.4 37.0 25.5
LLaVA-1.5-7B Liu et al. (2023a) 558K 665K 85.9 1510.7 316.1 64.3 58.3 66.1 30.5
LLaVA-1.5-13B Liu et al. (2023a) 558K 665K 85.9 1531.3 295.4 67.7 63.6 68.2 35.4
Multimodal Small Language Models
MobileVLM-3B Chu et al. (2023) 558K 158K 84.9 1288.9 - 59.6 - - -
MobileVLM-v2-3B Chu et al. (2024) 1.2M 2.4M 84.7 1440.5 - 63.2 - - -
MoE-LLaVA-3.6B Lin et al. (2024) 558K 1.59M 85.7 1431.3 - 68.0 - - 35.9
Bunny-3B He et al. (2024) 2M 695K 86.6 1488.8 289.3 68.6 - - -
Mipha-3B (Ours) 558K 665K 86.7 1488.9 295.0 69.7 64.3 68.9 32.1
[…]
Experimental Results on Instruction-Following Benchmarks. Furthermore, we evaluate our proposed Mipha on more challenging benchmarks that are designed for evaluation of MLLMs, i.e., POPE Li et al. (2023d), MME Fu et al. (2023), MMBench Liu et al. (2023d). MMBench-CN Liu et al. (2023d), SEED-Bench-img Li et al. (2023b), and MM-Vet Yu et al. (2023). As shown in Table 2, our proposed model, Mipha-3B, exceeds LLaVA-1.5-13B in performance on MMBench and MMBench-CN by 2% and 0.7%, respectively, despite both models being trained on the same dataset. Additionally, Mipha-3B demonstrates stronger resistance to hallucinations, surpassing LLaVA-1.5-13B on POPE by 0.8%. In comparison with other MSLMs, Bunny-3B, MoE-LLaVA-3.6B Lin et al. (2024), and MobileVLM-v2-3B Chu et al. (2024), we are astonished to find that despite Mipha-3B utilizing the least amount of training data, it achieved the best performance in 5 out of 6 benchmarks. In particular, on MMBench, Mipha-3B outperformed Bunny-3B by 1.1% and MoE-LLaVA by 1.7%, which are combined with the same visual backbone SigLIP and LM Phi-2-2.7B respectively. These results validate the effectiveness of our optimization strategies.
6 Conclusion
In this paper, we delve deeply into three aspects of Multimodal Small Language Models (MSLMs): the language models, visual representations, and optimization strategy. On the basis of our thorough analysis, we identify several improvements that can enhance the performance of MSLMs. Integrating insights from these aspects, we have developed a novel suite of models called Mipha. Compared to other MSLMs, our top-performing model, Mipha-3B, secures the highest performance in 9 out of 11 benchmarks. Moreover, when compared to the current state-of-the-art model, LLaVA-1.5-13B, Mipha-3B demonstrates comparable performance. We believe that Mipha offers a novel perspective on the training of strong MSLMs.

### 15 Mar 2024 v3

@@
 (2023b); Bai et al.
-(2023); Ye et al.
+(2023b); Ye et al.
 (2023a, b); Liu et al.
@@
 (2023e), Qwen-1.5 (1.8B) Bai et al.
-(2023), Pythia (available in both 1.4B and 2.8B sizes) Biderman et al.
+(2023b), Pythia (available in both 1.4B and 2.8B sizes) Biderman et al.
 (2023), and StableLM (1.6B) Stability (2024).
@@
 The first stage, or the pretraining stage, is dedicated to harmonizing vision and language features, establishing a solid foundation for multimodal understanding.
+Table 2: Multi-modal evaluation on 9 benchmarks: VQAv2 Goyal et al.
+(2017); GQA Hudson and Manning (2019); $\text{SQA}^{\text{I}}$: ScienceQA-IMG Lu et al.
+(2022); $\text{VQA}^{\text{T}}$: TextVQA Singh et al.
+(2019); MME-P: MME perception Fu et al.
+(2023); MME-C: MME cognition Fu et al.
+(2023); MMB: MMBench Liu et al.
+(2023d); MM-Vet Yu et al.
+(2023); POPE Li et al.
+(2023d).
+$*{}^{*}$he training images/annotations of the datasets are observed during training.
+${}^{\dagger}$ Includes in-house data that is not publicly accessible.
+Column Res.
+is the image resolution of visual backbone.
+Columns PT and IT are the data sizes in the pretraining stage and the visual instruction tuning stage, respectively.
+[…]
+Multimodal Large Language Models
+BLIP-2 Li et al. (2023c) Vicuna (13B) 224 129M - 65.0 41.0 61.0 42.5 1293.8 290.0 - 22.4 85.3
+InstructBLIP Dai et al. (2023b) Vicuna (7B) 224 129M 1.2M - 49.2 60.5 50.1 - - 36 26.2 -
+InstructBLIP Dai et al. (2023b) Vicuna (13B) 224 129M 1.2M - 49.5 63.1 50.7 1212.8 291.8 - 25.6 78.9
+Shikra Chen et al. (2023b) Vicuna (13B) 224 600K 5.5M $*{}^{*}$ 58.8 - -
+IDEFICS-9B Laurençon et al. (2023) LLaMA (7B) 224 353M 1M 50.9 38.4 - 25.9 - - 48.2 - -
+IDEFICS-80B Laurençon et al. (2023) LLaMA (65B) 224 353M 1M 60.0 45.2 - 30.9 - - 54.5 - -
+Qwen-VL Bai et al. (2023b) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 67.1 $*{}^{*}$ 38.2 - -
+Qwen-VL-Chat Bai et al. (2023b) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 68.2 $*{}^{*}$ 1487.5 360.71 60.6 - -
+mPLUG-Owl2 Ye et al. (2023b) LLaMA (7B) 448 400M 1.23M 79.4 56.1 68.7 58.2 1450.2 313.2 64.5 36.2 85.8
+LLaVA-1.5 Liu et al. (2023a) Vicuna (7B) 336 558K 665K $*{}^{*}$ $*{}^{*}$ 66.8 58.2 1510.7 316.1 64.3 30.5 85.9
+LLaVA-1.5 Liu et al. (2023a) Vicuna (13B) 336 558K 665K $*{}^{*}$ $*{}^{*}$ 71.6 61.3 1531.3 295.4 67.7 35.4 85.9
+Multimodal Small Language Models
+MobileVLM-1.7B Chu et al. (2023) M-LLaMA (1.4B) 336 558K 665K - 56.1 57.3 41.5 1196.2 - 53.2 - 84.5
+MobileVLM-3B Chu et al. (2023) M-LLaMA (2.7B) 336 558K 665K - 59.0 61.2 47.5 1288.9 - 59.6 - 84.9
+MobileVLM-v2-1.7B Chu et al. (2024) M-LLaMA (1.4B) 336 1.2M 2.4M - 59.3 66.7 52.1 1302.8 - 57.7 - 84.3
+MobileVLM-v2-3B Chu et al. (2024) M-LLaMA (2.7B) 336 1.2M 2.4M - 61.1 70.0 57.5 1440.5 - 63.2 - 84.7
+MoE-LLaVA-3.6B Lin et al. (2024) Phi-2 (2.7B) 384 558k 1.59M $*{}^{*}$ $*{}^{*}$ 70.3 57.0 1431.3 - 68.0 35.9 85.7
+Bunny-3B He et al. (2024) Phi-2 (2.7B) 384 2M 695K 79.8 62.5 70.9 - 1488.8 289.3 68.6 - 86.6
+Mipha-1.6B Phi-1.5 (1.3B) 384 558K 665K $*{}^{*}$ $*{}^{*}$ 58.3 45.6 1203.1 247.9 57.7 23.5 86.9
+Mipha-2.4B Gemma (2.0B) 384 558K 665K $*{}^{*}$ $*{}^{*}$ 65.3 52.4 1397.1 265.7 59.4 29.9 86.6
+Mipha-3B Phi-2 (2.7B) 384 558K 665K $*{}^{*}$ $*{}^{*}$ 70.9 56.6 1488.9 295.0 69.7 32.1 86.7
 The subsequent stage focuses on instructing the model to follow users’ instructions.
@@
 IDEFICS-80B Laurençon et al. (2023) LLaMA (65B) 224 353M 1M 60.0 45.2 36.0 - 30.9
-Qwen-VL Bai et al. (2023) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 35.2 67.1 $*{}^{*}$
-Qwen-VL-Chat Bai et al. (2023) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 38.9 68.2 $*{}^{*}$
+Qwen-VL Bai et al. (2023b) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 35.2 67.1 $*{}^{*}$
+Qwen-VL-Chat Bai et al. (2023b) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 38.9 68.2 $*{}^{*}$
 mPLUG-Owl2 Ye et al. (2023b) LLaMA (7B) 448 400M 1.23M 79.4 56.1 54.5 68.7 58.2
@@
 Multimodal Small Language Models
-MobileVLM Chu et al. (2023) M-LLaMA (2.7B) 336 558K 158K - 59.0 - 61.2 47.5
+MobileVLM Chu et al. (2023) M-LLaMA (2.7B) 336 558K 665K - 59.0 - 61.2 47.5
 MobileVLM-v2 Chu et al. (2024) M-LLaMA (2.7B) 336 1.2M 2.4M - 61.1 - 70.0 57.5
@@
 IDEFICS-80B Laurençon et al. (2023) 353M 1M - - - 54.5 38.1 53.2 -
-Qwen-VL Bai et al. (2023) ${}^{\dagger}$ ${}^{\dagger}$ 38.2 7.4 62.3 -
-Qwen-VL-Chat Bai et al. (2023) ${}^{\dagger}$ ${}^{\dagger}$ 1487.5 360.71 60.6 56.7 65.4 -
+Qwen-VL Bai et al. (2023b) ${}^{\dagger}$ ${}^{\dagger}$ 38.2 7.4 62.3 -
+Qwen-VL-Chat Bai et al. (2023b) ${}^{\dagger}$ ${}^{\dagger}$ 1487.5 360.71 60.6 56.7 65.4 -
 mPLUG-Owl2-7B Ye et al. (2023b) 400M 1.23M 85.8 1450.2 313.2 64.5 - - 36.2
@@
 Multimodal Small Language Models
-MobileVLM-3B Chu et al. (2023) 558K 158K 84.9 1288.9 - 59.6 - - -
+MobileVLM-3B Chu et al. (2023) 558K 665K 84.9 1288.9 - 59.6 - - -
 MobileVLM-v2-3B Chu et al. (2024) 1.2M 2.4M 84.7 1440.5 - 63.2 - - -
@@
 We believe that Mipha offers a novel perspective on the training of strong MSLMs.
+Appendix A Limitation
+While this work has extensively and thoroughly explored the design space of Multimodal Small Language Models (MSLMs), our analysis predominantly draws upon methodologies from current models, such as Phi-1.5/2 Li et al.
+[2023e], Javaheripi et al.
+[2023], StableLM Stability [2024], Pythia Biderman et al.
+[2023], and Gemma Google [2024], which are considered smaller language models.
+Given the rapid evolution of the field, the findings and conclusions of our study might not remain applicable to newer language models that emerge.
+Moreover, we recognize significant potential for further enhancements in Mipha, particularly in areas like Optical Character Recognition (OCR) and mathematical problem-solving.
+We plan to tackle these shortcomings in future research endeavors.
+Appendix B Negative Social Impact
+Despite Mipha showcasing advanced multimodal comprehension capabilities, it is not immune to the negative societal impacts observed in existing multimodal models.
+Foremost, Mipha has the potential to produce unreliable responses and may inadvertently perpetuate societal biases, including gender biases.
+Additionally, there exists the risk of Mipha being manipulated by users to generate harmful content.
+Appendix C Implementation Details
+The training process and dataset adopted by Mipha are substantially identical to LLaVA-1.5 Liu et al.
+[2023a].
+Specifically, Mipha adopts a two-stage training pipeline, i.e., the pretraining stage and the instruction tuning stage.
+Initially, we align the vision features from SigLIP Zhai et al.
+[2023] with the text embedding from Small Language Models (SLMs) Li et al.
+[2023e], Javaheripi et al.
+[2023].
+Subsequently, we apply visual instruction tuning to fully harness the capabilities of the Multimodal Small Language Models (MSLMs) across various multimodal tasks.
+We adopt the same cross-entropy loss during both stages for the next token prediction.
+In the pretraining stage, only the parameters of projector (a two-layer MLP) are updated on LCS-558K Liu et al.
+[2023a].
+The learning rate of 1e-3 and a batch size of 256 are adopted at this stage.
+In the instruction tuning stage, we train all parameters of Mipha for 2 epochs on a collection of mixture datasets with 665K samples, which is organized by Liu et al.
+[2023a].
+At this stage, we set the learning rate as 2e-5 and the batch size as 128.
+For both stages, we utilize the AdamW optimizer with no weight decay, characterized by momentum parameters of 0.9 and 0.98, and an epsilon value of 1e-7.
+We provide overall training settings in Fig 5.
+Additionally, for our model variants, we employ the same hyperparameters as Mipha, except for setting the epoch to 1 during the finetuning stage.
+We employ publicly available datasets to conduct instruct tuning for LMs.
+Specifically, we utilize the Alpaca dataset Taori et al.
+[2023] to supervised fine-tune Phi-2.
+For RHLF Ouyang et al.
+[2022] and Safe-RLHF Dai et al.
+[2023a], we employ the PKU-SafeRLHF-10K dataset Dai et al.
+[2023a] to finetune the models.
+Table 5: Hyper-parameters for training Mipha.
+Configuration Pretraining Instruction Tuning
+ViT init. SigLIP 384px/14 SigLIP 384px/14
+Projector init. Random Pretrain stage
+Image resolution $384\times 384$
+ViT sequence length 729
+LLM sequence length 2048
+Optimizer AdamW
+Optimizer hyperparameter $\beta_{1}=0.9,\beta_{2}=0.98,\epsilon=1e^{-7}$
+Peak learning rate $1e^{-4}$ $2e^{-5}$
+Training Epochs 1 2
+Global batch size 256 128
+Learning rate schedule Cosine
+Weight decay 0
+Warm-up ratio 0.03
+Numerical precision $\mathtt{bfloat16}$
+Activation checkpointing
+Figure 6: Additional Experimental Results: Selection of Small Language Models (Left) & Scaling Image Resolution (Right).
+C.1 More Experiments
+C.1.1 Selection of Small Language Models.
+Beyond the experiments presented in Fig.
+2 (left) of the main text, we have expanded our evaluation to include the performance of the Gemma-2B model Google [2024].
+We compared Gemma-2B with several smaller language models: Qwen-1.8B Bai et al.
+[2023a], Pythia-1.4B Biderman et al.
+[2023], StableLM-1.6B Stability [2024], and Phi-1.5-1.3B Li et al.
+[2023e].
+The results of these comparisons are illustrated in Fig 6 (left).
+The findings suggest that multimodal models leveraging Gemma-2B are versatile, achieving top-tier performance across the majority of benchmarks when compared to alternative methods.
+As one of the most recent additions to the small language model category, its performance on language-based tasks has exceeded that of LLaMA-2 Touvron et al.
+[2023b].
+We posit that Gemma-2B’s robust performance makes it an excellent foundational model for the development of MSLMs.
+C.1.2 Scaling Image Resolution.
+In addition to CLIP Radford et al.
+[2021], we also conduct experiments using SigLIP Zhai et al.
+[2023] with resolutions of 384px and 448px.
+Our findings, as illustrated inFig 6 (right), reveal that the SigLIP model with a 384px resolution outperforms its 448px counterpart in five out of eight benchmarks Hudson and Manning [2019], Singh et al.
+[2019], Goyal et al.
+[2017], Lu et al.
+[2022], Li et al.
+[2023d], Yu et al.
+[2023], Liu et al.
+[2023d], Fu et al.
+[2023].
+This suggests that simply increasing image resolution does not consistently improve the model’s performance across a diverse set of multimodal tasks.
+Additionally, with the exception of MM-Vet Yu et al.
+[2023], the two SigLIP variants show minimal variance in performance on the other seven benchmarks.
+This observation reinforces our assertion that higher image resolutions do not necessarily equate to better results in MSLMs.
+C.2 Qualitative Examples
+In Fig 7, 8 and 9, we present several cases from various perspectives such as commonsense reasoning, OCR, image description, object recognition, joke comprehension, and multilingual understanding.
+Mipha-3B demonstrates powerful detailed visual understanding by comprehensively answering questions based on the information in the images.
+Figure 7: Some cases of Commonsense Reasoning & OCR by Mipha-3B.
+For Commonsense Reasoning, Mipha-3B successfully identifies the Great Wall of Chain and the Great Pyramid of Giza in the image, going further to provide more detailed descriptions of them.
+Additionally, Mipha-3B also demonstrates strong OCR capabilities, such as successfully recognizing “since 1994” in the bottom right subfigure.
+RED means the critical part of the answer.
+Figure 8: Some cases of Image Description & Object Recognition by Mipha-3B.
+Mipha-3B is capable of inferring potential occurrences based on the scene, such as “a drag race event”, showcasing its robust visual perception abilities.
+Additionally, Mipha-3B demonstrates strong object recognition capabilities, including aspects such as counting, existence, and orientation.
+RED means the critical part of the answer.
+Figure 9: Some cases of Jokes Comprehension & Multilingual Understanding by Mipha-3B.
+Mipha-3B comprehends intriguing elements within the image and demonstrates multilingual abilities across Chinese and French.
+RED means the critical part of the answer.

### 25 Mar 2024 v4

@@
 Through detailed discussion, we provide insights and guidelines for developing strong MSLMs that rival the capabilities of MLLMs.
-Our code is available at https://github.com/zhuyiche/Mipha.
+Our code is available at https://github.com/zhuyiche/llava-phi.
 1 Introduction
@@
 MoE-LLaVA-3.6B Lin et al. (2024) Phi-2 (2.7B) 384 558k 1.59M $*{}^{*}$ $*{}^{*}$ 70.3 57.0 1431.3 - 68.0 35.9 85.7
-Bunny-3B He et al. (2024) Phi-2 (2.7B) 384 2M 695K 79.8 62.5 70.9 - 1488.8 289.3 68.6 - 86.6
+Bunny-3B He et al. (2024) Phi-2 (2.7B) 384 2M 695K 79.8 62.5 70.9 - 1488.8 289.3 68.6 - 86.8
 Mipha-1.6B Phi-1.5 (1.3B) 384 558K 665K $*{}^{*}$ $*{}^{*}$ 58.3 45.6 1203.1 247.9 57.7 23.5 86.9
@@
 We present and analyze four different configurations that were explored.
-Table 3: Multi-modal evaluation on academic-task-oriented benchmarks.
-VQAv2 Goyal et al.
-(2017); GQA Hudson and Manning (2019); VizWiz Gurari et al.
-(2018); $\text{SQA}^{\text{I}}$: ScienceQA-IMG Lu et al.
-(2022); $\text{VQA}^{\text{T}}$: TextVQA Singh et al.
-(2019).
-$*{}^{*}$he training images/annotations of the datasets are observed during training.
-${}^{\dagger}$ Includes in-house data that is not publicly accessible.
-Column Res.
-is the image resolution of visual backbone.
-Columns PT and IT are the data sizes in the pretraining stage and the visual instruction tuning stage, respectively.
-[…]
-Multimodal Large Language Models
-BLIP-2 Li et al. (2023c) Vicuna (13B) 224 129M - 65.0 41.0 19.6 61.0 42.5
-InstructBLIP Dai et al. (2023b) Vicuna (7B) 224 129M 1.2M - 49.2 34.5 60.5 50.1
-InstructBLIP Dai et al. (2023b) Vicuna (13B) 224 129M 1.2M - 49.5 33.4 63.1 50.7
-Shikra Chen et al. (2023b) Vicuna (13B) 224 600K 5.5M $*{}^{*}$
-IDEFICS-9B Laurençon et al. (2023) LLaMA (7B) 224 353M 1M 50.9 38.4 35.5 - 25.9
-IDEFICS-80B Laurençon et al. (2023) LLaMA (65B) 224 353M 1M 60.0 45.2 36.0 - 30.9
-Qwen-VL Bai et al. (2023b) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 35.2 67.1 $*{}^{*}$
-Qwen-VL-Chat Bai et al. (2023b) Qwen (7B) 448 ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$ $*{}^{*}$ 38.9 68.2 $*{}^{*}$
-mPLUG-Owl2 Ye et al. (2023b) LLaMA (7B) 448 400M 1.23M 79.4 56.1 54.5 68.7 58.2
-LLaVA-1.5 Liu et al. (2023a) Vicuna (7B) 336 558K 665K $*{}^{*}$ $*{}^{*}$ 50.0 66.8 58.2
-LLaVA-1.5 Liu et al. (2023a) Vicuna (13B) 336 558K 665K $*{}^{*}$ $*{}^{*}$ 53.6 71.6 61.3
-Multimodal Small Language Models
-MobileVLM Chu et al. (2023) M-LLaMA (2.7B) 336 558K 665K - 59.0 - 61.2 47.5
-MobileVLM-v2 Chu et al. (2024) M-LLaMA (2.7B) 336 1.2M 2.4M - 61.1 - 70.0 57.5
-MoE-LLaVA-3.6B Lin et al. (2024) Phi-2 (2.7B) 384 558k 1.59M $*{}^{*}$ $*{}^{*}$ 43.7 70.3 57.0
-Bunny-3B He et al. (2024) Phi-2 (2.7B) 384 2M 695K 79.8 62.5 - 70.9 -
-Mipha-3B (Ours) Phi-2 (2.7B) 384 558K 665K $*{}^{*}$ $*{}^{*}$ 47.5 70.9 56.6
   • Frozen ViT+SLM: In this configuration, the parameters of both the ViT and SLM are held constant, allowing the training to focus exclusively on aligning these two components via the projector.
@@
 5.2 Comparison with SOTA Models
-Table 4: Multi-modal evaluation on instruction following benchmarks.
-POPE Li et al.
-(2023d); MME-P: MME Perception Fu et al.
-(2023); MME-C: MME Cognition Fu et al.
-(2023); MMB: MMBench Liu et al.
-(2023d); SEED: SEED-Bench Li et al.
-(2023b); MM-Vet Yu et al.
-(2023).
-${}^{\dagger}$ Includes in-house data that is not publicly accessible.
-Columns PT and IT are the data sizes in the pretraining stage and the visual instruction tuning stage, respectively.
-Method PT IT POPE MME- MMB SEED MM-Vet
-P C en cn img
-Multimodal Large Language Models
-BLIP2-14B Li et al. (2023c) 129M - 85.3 1293.8 290.0 - - 49.7 22.4
-InstructBLIP-8B Dai et al. (2023b) 129M 1.2M - - - 36 23.7 58.8 26.2
-InstructBLIP-14B Dai et al. (2023b) 129M 1.2M 78.9 1212.8 291.8 - - - 25.6
-Shikra-13B Chen et al. (2023b) 600K 5.5M - - - 58.8 - - -
-IDEFICS-80B Laurençon et al. (2023) 353M 1M - - - 54.5 38.1 53.2 -
-Qwen-VL Bai et al. (2023b) ${}^{\dagger}$ ${}^{\dagger}$ 38.2 7.4 62.3 -
-Qwen-VL-Chat Bai et al. (2023b) ${}^{\dagger}$ ${}^{\dagger}$ 1487.5 360.71 60.6 56.7 65.4 -
-mPLUG-Owl2-7B Ye et al. (2023b) 400M 1.23M 85.8 1450.2 313.2 64.5 - - 36.2
-LLaVA-7B Liu et al. (2023b) 595K 158K 72.9 809.6 247.9 38.7 36.4 37.0 25.5
-LLaVA-1.5-7B Liu et al. (2023a) 558K 665K 85.9 1510.7 316.1 64.3 58.3 66.1 30.5
-LLaVA-1.5-13B Liu et al. (2023a) 558K 665K 85.9 1531.3 295.4 67.7 63.6 68.2 35.4
-Multimodal Small Language Models
-MobileVLM-3B Chu et al. (2023) 558K 665K 84.9 1288.9 - 59.6 - - -
-MobileVLM-v2-3B Chu et al. (2024) 1.2M 2.4M 84.7 1440.5 - 63.2 - - -
-MoE-LLaVA-3.6B Lin et al. (2024) 558K 1.59M 85.7 1431.3 - 68.0 - - 35.9
-Bunny-3B He et al. (2024) 2M 695K 86.6 1488.8 289.3 68.6 - - -
-Mipha-3B (Ours) 558K 665K 86.7 1488.9 295.0 69.7 64.3 68.9 32.1
 […]
-Experimental Results on Instruction-Following Benchmarks.
-Furthermore, we evaluate our proposed Mipha on more challenging benchmarks that are designed for evaluation of MLLMs, i.e., POPE Li et al.
-(2023d), MME Fu et al.
-(2023), MMBench Liu et al.
-(2023d).
-MMBench-CN Liu et al.
-(2023d), SEED-Bench-img Li et al.
-(2023b), and MM-Vet Yu et al.
-(2023).
-As shown in Table 3, our proposed model, Mipha-3B, exceeds LLaVA-1.5-13B in performance on MMBench and MMBench-CN by 2% and 0.7%, respectively, despite both models being trained on the same dataset.
-Additionally, Mipha-3B demonstrates stronger resistance to hallucinations, surpassing LLaVA-1.5-13B on POPE by 0.8%.
-In comparison with other MSLMs, Bunny-3B, MoE-LLaVA-3.6B Lin et al.
-(2024), and MobileVLM-v2-3B Chu et al.
-(2024), we are astonished to find that despite Mipha-3B utilizing the least amount of training data, it achieved the best performance in 5 out of 6 benchmarks.
-In particular, on MMBench, Mipha-3B outperformed Bunny-3B by 1.1% and MoE-LLaVA by 1.7%, which are combined with the same visual backbone SigLIP and LM Phi-2-2.7B respectively.
-These results validate the effectiveness of our optimization strategies.
 6 Conclusion
@@
 Integrating insights from these aspects, we have developed a novel suite of models called Mipha.
-Compared to other MSLMs, our top-performing model, Mipha-3B, secures the highest performance in 9 out of 11 benchmarks.
+Compared to other MSLMs, our top-performing model, Mipha-3B, secures the highest performance in 8 out of 11 benchmarks.
 Moreover, when compared to the current state-of-the-art model, LLaVA-1.5-13B, Mipha-3B demonstrates comparable performance.
@@
 This observation reinforces our assertion that higher image resolutions do not necessarily equate to better results in MSLMs.
+C.1.3 More Experimental Result on Mipha.
+Table 4: Multi-modal evaluation on instruction following benchmarks.
+POPE Li et al.
+[2023d]; MME-P: MME Perception Fu et al.
+[2023]; MME-C: MME Cognition Fu et al.
+[2023]; MMB: MMBench Liu et al.
+[2023d]; SEED: SEED-Bench Li et al.
+[2023b]; MM-Vet Yu et al.
+[2023].
+${}^{\dagger}$ Includes in-house data that is not publicly accessible.
+Columns PT and IT are the data sizes in the pretraining stage and the visual instruction tuning stage, respectively.
+Method PT IT POPE MME- MMB SEED MM-Vet
+P C en cn img
+Multimodal Large Language Models
+BLIP2-14B Li et al. [2023c] 129M - 85.3 1293.8 290.0 - - 49.7 22.4
+InstructBLIP-8B Dai et al. [2023b] 129M 1.2M - - - 36 23.7 58.8 26.2
+InstructBLIP-14B Dai et al. [2023b] 129M 1.2M 78.9 1212.8 291.8 - - - 25.6
+Shikra-13B Chen et al. [2023b] 600K 5.5M - - - 58.8 - - -
+IDEFICS-80B Laurençon et al. [2023] 353M 1M - - - 54.5 38.1 53.2 -
+Qwen-VL Bai et al. [2023b] ${}^{\dagger}$ ${}^{\dagger}$ 38.2 7.4 62.3 -
+Qwen-VL-Chat Bai et al. [2023b] ${}^{\dagger}$ ${}^{\dagger}$ 1487.5 360.71 60.6 56.7 65.4 -
+mPLUG-Owl2-7B Ye et al. [2023b] 400M 1.23M 85.8 1450.2 313.2 64.5 - - 36.2
+LLaVA-7B Liu et al. [2023b] 595K 158K 72.9 809.6 247.9 38.7 36.4 37.0 25.5
+LLaVA-1.5-7B Liu et al. [2023a] 558K 665K 85.9 1510.7 316.1 64.3 58.3 66.1 30.5
+LLaVA-1.5-13B Liu et al. [2023a] 558K 665K 85.9 1531.3 295.4 67.7 63.6 68.2 35.4
+Multimodal Small Language Models
+MobileVLM-3B Chu et al. [2023] 558K 665K 84.9 1288.9 - 59.6 - - -
+MobileVLM-v2-3B Chu et al. [2024] 1.2M 2.4M 84.7 1440.5 - 63.2 - - -
+MoE-LLaVA-3.6B Lin et al. [2024] 558K 1.59M 85.7 1431.3 - 68.0 - - 35.9
+Bunny-3B He et al. [2024] 2M 695K 86.8 1488.8 289.3 68.6 - - -
+Mipha-3B (Ours) 558K 665K 86.7 1488.9 295.0 69.7 64.3 68.9 32.1
+[…]
+Experimental Results on Instruction-Following Benchmarks.
+Furthermore, we evaluate our proposed Mipha on more challenging benchmarks that are designed for evaluation of MLLMs, i.e., POPE Li et al.
+[2023d], MME Fu et al.
+[2023], MMBench Liu et al.
+[2023d].
+MMBench-CN Liu et al.
+[2023d], SEED-Bench-img Li et al.
+[2023b], and MM-Vet Yu et al.
+[2023].
+As shown in Table LABEL:tbl:qa_table, our proposed model, Mipha-3B, exceeds LLaVA-1.5-13B in performance on MMBench and MMBench-CN by 2% and 0.7%, respectively, despite both models being trained on the same dataset.
+Additionally, Mipha-3B demonstrates stronger resistance to hallucinations, surpassing LLaVA-1.5-13B on POPE by 0.8%.
+In comparison with other MSLMs, Bunny-3B, MoE-LLaVA-3.6B Lin et al.
+[2024], and MobileVLM-v2-3B Chu et al.
+[2024], we are astonished to find that despite Mipha-3B utilizing the least amount of training data, it achieved the best performance in 4 out of 6 benchmarks.
+In particular, on MMBench, Mipha-3B outperformed Bunny-3B by 1.1% and MoE-LLaVA by 1.7%, which are combined with the same visual backbone SigLIP and LM Phi-2-2.7B respectively.
+These results validate the effectiveness of our optimization strategies.
 C.2 Qualitative Examples